# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

### v3

@@ 8 Concluding remarks
This paper introduces three preconditioners, termed GSS, RGSS-I and RGSS-II for solving DSPPs that arise in various applications. We provide a convergence analysis for the GSS iterative method, demonstrating that the method converges for any initial guess vector when the parameter $\bm{\omega}\geq 1/2.$ Moreover, spectral bounds for the preconditioned matrices are derived. Additionally, we have shown that the RGSS-II preconditioner requires almost $m+1$ iterations to solve the DSPP. Numerical experiments for the DSPP arising from the PDE-constrained optimization problem are performed, which demonstrate that the proposed preconditioners are efficient and outperform the existing state-of-the-art preconditioners.
@@ Acknowledgments
Pinki Khatun gratefully acknowledges the Council of Scientific &\&& Industrial Research (CSIR), New Delhi, India, for providing financial support through a fellowship (File No. 09/1022(0098)/2020-EMR-I).

### v4

@@ 8 Concluding remarks
This paper introduces three preconditioners, termed GSS, RGSS-I and RGSS-II for solving DSPPs that arise in various applications. We provide a convergence analysis for the GSS iterative method, demonstrating that the method converges for any initial guess vector when the parameter $\bm{\omega}\geq 1/2.$ Moreover, spectral bounds for the preconditioned matrices are derived. Additionally, we have shown that the RGSS-II preconditioner requires at most $m+1$ iterations to solve the DSPP. Numerical experiments for the DSPP arising from the PDE-constrained optimization problem and leaky lid driven cavity problem are performed, which demonstrate that the proposed preconditioners are efficient and outperform the existing state-of-the-art preconditioners. A line for future research direction is to investigate the spectral bounds of the RGSS-I and RGSS-II preconditioners when the diagonal blocks are nonsymmetric.
@@ Acknowledgments
The authors sincerely acknowledge the anonymous reviewers for their constructive feedback to improve the paper. Pinki Khatun gratefully acknowledges the Council of Scientific &\&& Industrial Research (CSIR), New Delhi, India, for providing financial support through a fellowship (File No. 09/1022(0098)/2020-EMR-I).